OpenStudy (anonymous):

Write the equation of the line that passes through (3, 6) and (0, –3) in slope-intercept form.

OpenStudy (anonymous):

find the slope m = (y2-y1)/(x2-x1) then put it here y=mx+c substitute the value of m